### v1

# Bar Evolution in Edge-on Galaxies: A Demographic Study of Boxy/Peanut Bulges

Abstract
Context. Boxy/peanut and X-shaped (BP/X) bulges are prominent features in edge-on disk galaxies and are believed to be vertically thickened bars. Despite their relevance in bar evolution, a statistically robust census of these structures in large surveys has been lacking.
Aims. We aim to provide the largest catalog of BP/X structures in edge-on galaxies to date, and to investigate their properties and role in shaping galaxy scaling relations.
Methods. We selected a sample of 6684 edge-on galaxies from SDSS DR8 using Galaxy Zoo classifications, requiring a high edge-on probability () and a minimum of 10 independent votes. Two-dimensional image decomposition is performed using GALFIT to obtain structural parameters. Residual images are visually inspected to classify BP/X features into four categories: strong both-sided, both-sided, one-sided, and control (no BP/X). We also estimated stellar mass, distance, and physical size for each galaxy.
Results. Out of 6653 classified galaxies, we identified 1675 (25%) with both-sided BP/X features—545 (8%) strong and 1130 (17%) faint—as well as 1108 (17%) one-sided structures, making up a total of 2783 BP/X-hosting galaxies (42%). One-sided structures, likely signatures of ongoing buckling, are more frequent than strong both-sided bulges across all stellar masses. The fraction of BP/X bulges increases with stellar surface mass density, indicating a connection with bar formation in dense disks. We also find that galaxies with strong BP/X bulges contribute to increased scatter in the stellar mass–size and stellar mass–surface density relations, particularly at higher masses.
Key Words.:
Galaxies: formation – Galaxies: evolution1 Introduction
Spiral galaxies with elongated bar-like structures in their center are commonly known as barred galaxies and comprise nearly 30-70% of all spiral galaxies (Aguerri et al., 2009; Masters et al., 2011; D´ıaz-Garc´ıa et al., 2016). Bars play a major role in the evolution of spiral galaxies (Marchuk et al., 2022). They are thought to funnel a large amount of gas into the nuclear region (Shlosman et al., 1989; Boone et al., 2007). This gas helps in the formation of new stars in the central and bar regions (D´ıaz-Garc´ıa et al., 2020; Fraser-McKelvie et al., 2020). The accretion of gas into the nuclear region also feeds the central supermassive black holes (SMBHs) that help the growth of SMBHs. Sometimes the excess gas accretion onto SMBH ignites the nuclear activity in the galaxy (Alonso et al., 2013; Galloway et al., 2015; Silva-Lima et al., 2022; Kataria & Vivek, 2024).
During the growth phase of bars, nuclear stellar disc (NSD) is formed due to intense star formation in the central sub-kpc region (Friedli & Benz, 1993; Heller & Shlosman, 1994; Friedli & Benz, 1995; Martin & Friedli, 1997; Wozniak, 2007; Kim et al., 2011; Cole et al., 2014; Seo et al., 2019; Baba & Kawata, 2020; Sormani et al., 2022). The bar suppresses star formation through out the remainder of its extent (Martin & Friedli, 1997; Spinoso et al., 2016; Khoperskov et al., 2018; Donohoe-Keyes et al., 2019). Therefore, the NSD has younger stars than the age of the bar (Baba & Kawata, 2020) and the BPX bulge contains stars older than the NSD (Gonzalez et al., 2017). Using their numerical simulations, Baba & Kawata (2020) showed that the oldest stellar population in the NSD can be used for dating bar formation time (also see de Sá-Freitas et al., 2023).
Bars drive the secular evolution of spiral galaxies by redistributing matter and angular momentum (Athanassoula, 2003; Minchev et al., 2011). Despite numerous studies on bar formation and evolution, their origin remains a topic of debate (see Sellwood, 2014, for a review). The bar formation mechanisms involve global unstable modes of rotating disk (Kalnajs, 1972; Ostriker & Peebles, 1973), swing amplification of leading small perturbations (Toomre, 1981), slow trapping of stellar orbits due to a weak non-axisymmetric perturbation (Lynden-Bell, 1979), mutual gravitational interaction of precessing orbits (Polyachenko & Polyachenko, 2003), and tidal interactions between galaxies (Miwa & Noguchi, 1998; Lang et al., 2014; Łokas, 2018).
Studies suggest that disc evolution of the galaxy also includes the development of bars in length and thickness (Athanassoula, 2003, 2005). The vertical thickening of bars in galaxies is mainly studied using numerical simulations (Combes et al., 1990; Raha et al., 1991; Bureau & Athanassoula, 2005; Martinez-Valpuesta et al., 2006; Kumar et al., 2021, 2022; Kumar, 2023; Ghosh et al., 2024). Among all bar thickening mechanisms bar buckling is the most violent mechanism, in which the galactic bar bends out of the galactic plane (Combes et al., 1990; Raha et al., 1991; Łokas, 2019; Kumar et al., 2022). Buckling instability involves spontaneous symmetry breaking with respect to the galactic equatorial plane (Pfenniger & Friedli, 1991; Raha et al., 1991), resulting in the thickening and weakening of bar over dynamical timescales. Observation of some galaxies having Boxy/Peanut bulge shows the presence of a bar at their center, suggesting that B/P bulge is a by-product of galactic bar formation (Bettoni & Galletta, 1994; Kuijken & Merrifield, 1995; Veilleux et al., 1999).
Alternative mechanisms proposed to explain the development of BPX-shaped bulges include vertical resonant heating (Combes et al., 1990; Pfenniger & Friedli, 1991; Quillen et al., 2014) or resonant trapping within a vILR (vertical inner Lindblad resonance) during the secular evolution of bars (Quillen, 2002; Sellwood & Gerhard, 2020). Recently, Zozulia et al. (2024a, b, 2025) investigated bar formation and thickening through the analysis of action variables and resonant angles of stellar orbits. Their results demonstrate that vertical trapping plays a dominant role in shaping the boxy/peanut morphology of galactic bars. Recent N-body simulations (Smirnov & Sotnikova, 2018; Sellwood & Gerhard, 2020) have illustrated that these mechanisms can engender BP/X-shaped bulges without immediate bar buckling following bar formation.
We observe the formation of boxy/peanut (BP/X) bulges in a wide range of numerical simulations (e.g., Friedli & Pfenniger, 1990; Raha et al., 1991; Merritt & Sellwood, 1994; Debattista et al., 2006; Martinez-Valpuesta et al., 2006; Saha et al., 2013; Fragkoudi et al., 2017; Debattista, 2018; Smirnov & Sotnikova, 2018; Di Matteo et al., 2019; Khoperskov et al., 2019; Łokas, 2019; Smirnov & Sotnikova, 2019; Collier, 2020; Debattista et al., 2020; Kumar et al., 2022). Prior to bar formation, galaxies typically exhibit vertical symmetry. However, three-dimensional models of bar-unstable disks reveal that newly formed bars can bend out of the disk plane, losing coherence and inducing random vertical motions (Raha et al., 1991). This leads to a bar that is significantly thicker than the surrounding disk and appears boxy or peanut-shaped when viewed edge-on (Combes & Sanders, 1981; Combes et al., 1990). Such bulges are now widely interpreted as edge-on bars. The vertical asymmetry arises after the bar formation and grows rapidly indicating a clear peak followed by rapid decreases within a few gigayears (Athanassoula & Martinez-Valpuesta, 2008; Łokas, 2019). Recently, Erwin & Debattista (2016); Li et al. (2017); Xiang et al. (2021) reported evidence of buckling instability in total eight observed galaxies, providing support for buckling driven BP/X bulges.
There have been a few attempts to study the set of galaxies having BP/X bulges (Ciambur & Graham, 2016; Laurikainen & Salo, 2016; Savchenko et al., 2017). However, it is not easy to detect the presence of a BP/X bulge, as it requires the bulge to be oriented in a perpendicular direction with respect to our line of sight. There has been an attempt to create a catalogue of edge-on galaxies (Kautsch et al., 2006). Additionally, there have been attempts to study the relationship between the development of BP/X bulges and the dark matter halo using numerical simulations (Kumar et al., 2022). Several previous studies (discussed in Section 7) have shown a strong correlation between stellar mass and the presence of BP/X bulges. These studies have been conducted with relatively smaller sample sizes than ours. They concluded that high-mass galaxies are more likely to show the presence of a BP/X bulge than lower-mass galaxies. The gas mass fraction could play a role in determining the way in which the bar develops—either via buckling or symmetric growth—as suggested by Berentzen et al. (2007). However, the current gas mass fraction does not appear to influence whether a BP/X bulge will form (Erwin & Debattista, 2017).
Here, we aim to separate the edge-on galaxies having BP/X bulge from a large sample of edge-on galaxies obtained from the Sloan Digital Sky Survey111SDSS. We classify galaxies having BP/X bulges and study different parameters like masses and sizes of galaxies with or without BP/X bulges. Then we compare their properties to find any speciality in the sample with BP/X bulges. We verified some of the past similar analysis performed previously on relatively smaller sample size. Further, we investigate the effect of BP/X hosting galaxies on the scaling laws, mass-size relation.
2 Sample Selection
2.1 The data base
As a first step of our study, we aim to compile a set of edge-on galaxies as BP/X bulges are only identifiable in edge-on projection when bar is nearly side-on to our line-of-sight. For this purpose, we obtain the ZooSpec table from the Galaxy Zoo 1 project (Lintott et al., 2008, 2010) available with SDSS data. The Galaxy Zoo is a citizen science program where volunteer citizens help astronomers identify several unique morphological features, such as bars, spiral arms, bulge, rings etc. by voting on an interactive web interface. The ZooSpec table provide us with the voting data which we use to select the set of edge-on galaxies for our analysis. Followings are the two criteria we have imposed to select desired galaxies:
1.
number of votes (nvote) ,
2.
fraction of votes for edge-on galaxy (p_edge) ¿ 0.9
The main motivation for choosing minimum 10 votes for each galaxy is to avoid an uncertain classification due to small number of voting. A fraction greater than 0.9 signifies that the most volunteers have identified the galaxy as edge-on. Then we obtain coordinates (RA, Dec) of the selected galaxies from the ”Galaxy” table of the Sloan Digital Sky Survey data release 8 (SDSS DR8)222SDSS DR8. The ”Galaxy” table contains photometric data of all the objects classified as galaxies from the ”PhotoPrimary” table. Though SDSS DR17 is currently available, we use DR8 because ZooSpec table builds on the DR7 and its link is available with DR8 (Aihara et al., 2011). These conditions return us a set of 9068 edge-on galaxies. The SQL search query for this task is attached below.
SELECT g.ra, g.dec, g.sky_r, g.petroRad_r, g.sky_r, g.expPhi_r FROM Galaxy g JOIN zooSpec z ON z.objid = g.objID WHERE z.nvote > 10 AND z.p_edge > 0.90
2.2 FITS image files
Next, we use the (RA, Dec) of the egde-on galaxies to download the r-band images in FITS (Flexible Image Transport System) format. Our choice of -band images reflects our interest in the stellar light profile of galaxies because -band is less affected by dust extinction and has higher S/N ratio (York et al., 2000), which makes it easier to distinguish other structures from the oval (Yoshino & Yamauchi, 2015). For this purpose, we utilise the module 333Skyview of the python package Astroquery (Ginsburg et al., 2019). Astroquery provide a set of tools for accessing astronomical data from the webs of various surveys. To obtain the complete image of the galaxy, we use the radius of the field to be 1.5 times the Petrosian radius of the galaxy. The pre-computed Petrosian radii of the galaxies are obtained using SDSS SQL search query along with the (RA, Dec) positions as discussed in Section 2.1. For all the galaxies, we used the default Skyview image dimensions, i.e., 300 pixels 300 pixels. Due to our default image dimensions, the pixel size varies among galaxies depending upon the corresponding petrosian radius. Out of 9068 galaxies, we can obtain images for 6684 galaxies because Skyview does not return images for some specific positions. Through a quick visual inspection we removed incomplete images (where only some part of the galaxy is visible) leaving a sample of 6653 galaxies for further analysis. For all the galaxies in our sample we download the corresponding PSF file from SDSS Science Archive Server.
2.3 Other relevant parameters
Along with obtaining FITS images of the edge-on galaxies, we also complied some relevant parameters which we will use later in this study. These parameters include object ID (objid), redshift (z), absolute magnitudes. For this purpose we make use of (RA, DEC) positions of selected galaxies in SDSS crossID tool 444SDSS crossID. This tool allows users to upload their own table which can be joined with different SDSS tables to extract other information using a similar SQL query as described in Section 2.1. We extract the object ID (objid), absolute magnitude in -, -, & -band from Photoz table, and redshift () from the SpecObjAll table. We also obtain exponential fitting parameters like minor to major axis ratio (), effective radius (Re), position angle, and magnitude from the PhotoObjAll table.
3 FITS images processing
As outlined earlier in the introduction, this work focuses on the characteristics of BP/X bulges and their host galaxies, we tried some well-known methods (e.g., median filtering, ellipse fitting, profile fitting, etc.) to identify the BP/X structures in our sample. In this section, we discuss the methods we tried, their drawbacks, and finally adopted method for further analysis.
3.1 Median/Gaussian filter
Using the median or Gaussian filter to smooth the image, and then subtracting these smoothed image from original image (and sometimes adding a factor of residual image back into the original image) to sharp the hidden BP/X structure has been used in the literature (Faúndez-Abans & de Oliveira-Abans, 1998). Therefore, we first started with using the median filter available in the Scipy Python package555Scipy (Virtanen et al., 2020a) to identify the BP/X hosting galaxies. The median filter smooth any image according to input box (or window) size. We give galaxy image and the box size as input parameters. The value of each pixel in the output image is calculated as the median value of the pixels inside the box of the given size around that pixel. Then we subtract the median filtered image from the original image to get the residual image. The residual image shows the salient features hidden near the central part of the galaxy as the outer regions mostly remain the same in both images and get cancel out in the residual image.
We tested this method on some images and found that the clarity of the structure in the residual image varies with the box size used for smoothing. For each image, we need to find a suitable box size by trial and error method. We could not find any specific box size that can work for all images. This method provides a work around when we want to search BP/X structures in a few images. Otherwise, it is way difficult to manually identify a correct box size for individual image given large sample like we have in this study. Therefore, we do not use median/Gaussian filter in our analysis.
3.2 Ellipse fitting
The subtraction of smooth intensity distribution of galaxy from its observed intensity distribution can also sharpen the hidden BP/X structure. Therefore, we construct smooth intensity distribution of galaxy by fitting iso-intensity ellipses using a Python package Photutils 666ellipse.photutils (Bradley & Developers (2023)). It requires an initial guess for an ellipse geometry to start the fitting. The initial ellipse is described by the user guess for center, semi-major axis, ellipticity, and position angle. We guess initial parameters to perform ellipse fitting, and use the fitted ellipses to model the smooth intensity profile of the galaxy. Then we subtract this model image from observed image to find the residual image. This residual images show the hidden structures behind the smooth intensity distribution of galaxy.
To automate this process for our large sample, we write a Python script to calculate the center of mass and perform Fourier decomposition of each image. The center of mass goes as center of ellipse, phase angle of Fourier mode goes as position angle, and the radius where Fourier mode reaches maximum goes as semi-major axis. For ellipticity (), we use minor to major axis ratio () obtained from Galaxy table of SDSS data, i.e., . Note that we would have used initial guesses (e.g., position angle and radius) from exponential fitting obtained from SDSS data for ellipse fitting (Stoughton et al., 2002). However, we performed Fourier analysis to minimize the failure of automatic fitting process because SDSS parameters are very different for some galaxies, e.g., some SDSS edge-on galaxies do not look like edge-on. This method works quite well as compare to median/Gaussian filter method for large sample. However, it fails fitting frequently when there is some discontinuity in the image or guessed parameters are somewhat different. For example, presence of other galaxy, bright foreground star, dust lane, etc.
3.3 galfit profile fitting
Another widely used method to enhance the visibility of hidden structures is to subtract a smooth model intensity distribution from observed intensity distribution (Yoshino & Yamauchi, 2015), (Ness & Lang, 2016). Since most of the disk galaxies follow exponential intensity distribution, we can model exponential intensity distribution using exponential fitting parameters obtained from SDSS data. However, as mentioned earlier in Section 3.2, not all the edge-on SDSS galaxies show edge-on intensity distribution. Also, a radial exponential profile used in SDSS pipeline (Stoughton et al., 2002) does not completely describe disk profile (van der Kruit & Freeman, 2011). Therefore, we do not directly use the SDSS exponential fitting parameters. Instead, we performed 2D fitting using edge-on disk profile available in galfit 2D image decomposition tool (Peng et al., 2010). The edge-on disk profile is expressed in terms of scale radius (Rs) and scale height () as follows (Van der Kruit & Searle, 1981),
where is the central surface brightness, is a Bessel function.
galfit is a stand-alone 2-D image decomposition program written in C language. It uses an arbitrary number and mix of parametric functions (such as Sersic, Nuker, Gaussian, Moffat, and exponential) to decompose the image. It creates the model image and iterates it over and over the original image to find the perfect fit. It is capable of fitting galaxies and the sub-structures in them. Although the presence of dust significantly affects photometric decomposition in edge-on galaxies, its impact must be carefully considered. Using radiative transfer modeling of simulated galaxies, Savchenko et al. (2023) showed that dust can lead to a 25–50% increase in the fitted disk scale radius in edge-on systems. In contrast, we expect bar formation to enhance central brightness, which typically results in a reduced disk scale radius in single component fitting, while bar thickening is expected to produce counter effect similar to presence of the dust. A detailed investigation of the counterbalancing effects of dust and bar presence is beyond the scope of this paper. In the case of extensive data like in this study, automation is very essential. For automation, user need to provide an external ‘wrapping’ algorithm for taking care of the pre-processing and post-processing of the fittings. For our image processing, we provided an input file containing all the input parameters (initial guesses) to start the computation. Next, we discuss these initial guesses.
Initial guesses
To initialize the fitting process in galfit we need to provide initial guesses for central surface brightness, scale radius and scale height in equation (1) along with center and position angle. We obtain the initial guesses for center, position angle and scale radius using center of mass and Fourier analysis as discussed in Section 3.2. Being the sample of edge-on disk galaxies, the initial guess for scale height is assumed to be 0.2 scale radius. If fitting fails, we try with larger scale height of 0.5 scale radius, particularly when galaxies do not look like edge-on. Since galfit performs edge-on disk surface brightness fit in . To guess the central surface brightness, we consider a square of 10 pixels 10 pixels size around the center of galaxy, and measure total counts in this box as ADUs. Then, we measure the central surface brightness in using following expression,
where and are the side lengths of the region in x and y direction respectively, here taken as 10 pixels in each directions. Time is the exposure time, which is taken to be 53.9 seconds for the r-band. Magnitude is the zero point magnitude, here taken to be 24.8 (see Table 21 of Stoughton et al., 2002).
For each galaxy, all these initial guess parameters are compiled into a text file in galfit input format. It also includes galaxy name, output file name, convolution box size among other parameters. Our wrapping algorithm for automation takes care of the variable parameters, e.g., initial guesses, galaxy name, output file name, etc. This fitting analysis provides us with edge-disk parameters which are not available in the original SDSS catalogue. In Fig. 1, we show comparison of exponential disk parameters obtained from SDSS and edge-disk parameters from our galfit decomposition. The left panel compares position angles, and the middle panel compares scale radii. Our analysis shows a good agreement between SDSS and galfit position angles and scale radii as evident from the equality line displayed in left most and middle panels. The SDSS exponential fitting semi-minor to semi-major axis ratio () and galfit edge-disk fitting scale height to scale radius ratio () exhibits linear relation but non-unity slope as shown in the right most panel. We fit the straight line and obtain following relation,
4 BP/X bulge Classification
After edge-disk fitting using galfit, we visually inspected residual images ( original image model image) of all 6653 galaxies to build a catalogue of galaxies with and without BP/X bulges for detailed study. The sample without BP/X bulges is considered as control sample for comparison with the sample having BP/X bulges. We identify BP/X bulges in different stages/contrast. Following is our classification scheme for BP/X bulges and control sample.
Both-sided:- This class of galaxies show the presence of x-shape or peanut-shape structure at their center in the residual images. The structure can be clearly visible as in Fig. 2(a,b,c) or can be faintly visible as in Fig. 2(d,e,f). We identify 1675 galaxies in this class, which comprise nearly of the total sample.
Strong both-sided:- From the class of both-sided BP/X bulge galaxies, we separate galaxies showing a clear x-shape structures and classify them into strong both-sided sample. These galaxies show a very prominent x-shape structure as in Fig. 2(a,b,c). There are 545 galaxies (nearly ) in this class.
One-sided:- The residual images of these galaxies shows excess intensity on one side of the galactic mid-plane. It seems like the bulge is present on only one side of the galaxy. Fig. 2(g,h,i) is an example of the galaxy in this sample. One of the possible reason for this appearance is the dust obscuration on one side. Additionally, it is also possible that this class potentially contains galaxies in the initial buckling phase. There are 1108 galaxies in this class, which is nearly of the total sample.
Control:- These galaxies show no sign of BP/X bulge formation. Therefore, the model completely fits the original image, leaving no feature in the residual. There are a total of 3870 galaxies in this class. Some of the galaxies in the control sample show presence of other significantly bright structure in the projection, e.g., in the bottom left side of Fig. 2(j,k,l). There are 146 galaxies which show nearby bright structures (likely neighbor galaxy). The remaining 3724 galaxies do not show any neighboring bright objects in the projection. So, nearly of our sample belongs to this class. Note that some of the galaxies showing showing spheroidal structures (classical bulge or end-on projected bars) are also included in the control sample.
5 Derived Quantities for Galaxies
Along with the parameters obtained from SDSS catalogue and measured using 2D image decomposition using galfit, we derive stellar masses, distances and physical sizes of the galaxies from existing information. The following subsections describes our methodology for measuring these derived properties of galaxies.
5.1 Stellar masses
To proceed with the stellar mass estimation, we require a mass-to-light ratio relation. For this purpose, we use a widely adopted stellar mass-to-light ratio relation available for SDSS filters.
where is the mass-to-light ratio of the galaxy in imaging band. Parameters and are two constants for the same band at a given color index (here, ). We have obtained these constants from Bell et al. (2003); Du & McGaugh (2020). Both, Bell et al. (2003) and Du & McGaugh (2020) report similar and parameters for the color. Therefore, we consider using and magnitudes for the stellar mass estimation of our sample galaxies (Kumar & Kataria, 2022). We obtain the galactic extinction corrected - and - band magnitude from SDSS PhotoObjAll table. The internal extinction in - band is calculated using the relation , where is obtained from Shao et al. (2007). Using the relation (see Table 2 of Yuan et al., 2013), we measure the internal extinction in -band. Then the internal extinction corrected color and magnitude are given as
We use the extinction corrected absolute magnitudes and colors for mass calculation.
After calculating mass-to-light ratio using above described method, we use absolute magnitudes (Mr) to measure the luminosities of galaxies in solar unit by following expression.
where, M⊙ is solar magnitude in . For our analysis, we obtained solar magnitude from Blanton et al. (2003). Next, we measure stellar mass of the galaxies in solar mass unit by multiplying mass-to-light ratio with their luminosities as follows,
5.2 Angular distances and physical sizes
We use the spectroscopic redshifts of galaxies obtained from SDSS database to measure their angular distances. In the flat universe, the angular distance of an object at redshift is given by the following expression (Simard et al., 2011).
where is the speed of light in vacuum, is the Hubble’s constant, is the matter density of the universe, and is the dark energy density. For this study, we have considered , , and . The peculiar velocities of nearby galaxies affect the redshift measurement, and thus their distance using above expression. Nonetheless, all the galaxy in our sample have , indicating marginal effect of peculiar velocities on the distance estimation (Shen et al., 2003).
From the galfit photometric decomposition, we obtain the disk scale height and disk scale radius in pixels. Since we have downloaded the FITS images of galaxies within 3 Petrosian radius aperture with 300 pixels 300 pixels size, the pixel size is simply Petrosian radius / 100 in the same unit as Petrosian radius (here, ). Using the pixel size we calculate the angular size of galaxies in radian. Then we obtained the physical sizes of galaxies by multiplying angular sizes and angular distances.
6 Sample Characteristics
As mentioned in Section 2.2, we started with 6684 edge-on galaxies, and performed 2D morphological decomposition using galfit. Further, we visually separated these objects into four different subclasses depending on central features as discussed in Section 4. Fig 3 shows the distributions of magnitudes, redshifts, scale radii, and stellar masses of our complete sample. The median values of the distributions are represented with the red color vertical lines at the bottom axes of respective panels. The lower and upper limits of dark distributions respectively indicates and range as shown in the legends of respective panels.
Top left panel of Fig. 3 shows the absolute magnitude distribution of the sample galaxies. Our sample cover an absolute band magnitude from to . The median of magnitude distribution lies at with percentile range from -18.6 to -21.0. As most of the galaxies are bright and resolved, they are primarily local volume objects. The median redshift of the sample is 0.06 with percentile range from 0.03 to 0.09 as can be seen from top right panel of the figure. The farthest galaxy in the sample lies at redshift 0.2.
The bottom left panel of Fig. 3 displays the disk scale radius of galaxies. The median angular scale radius of the sample is with range in band. A few nearby galaxies show larger angular size greater than . For clarity, we have shown the distribution up to . Our derived stellar masses of the galaxies span a wide range from to . The median stellar mass of the sample is and the percentile range covers to (see the bottom right panel of Fig. 3).
7 Result
7.1 Mass and size distributions of BP/X bulges
First, we investigate the differences in properties like mass, disk radius, disk height and axial ratio among the different classes we have identified in the whole sample. Fig. 4, we plot the distributions of mass, axis ratio, disk height, and disk radius for for samples in this study. The control, both-sided, one-sided, and strong both-sided are respectively shown with blue, green, red, and purple colors. To better understand the distribution, we fitted each dataset using the ’log normal’ distribution. For fitting, we used the Python package scipy .777scipy
The distribution of mass (4a) shows that the mass distribution for the samples without any signature of bulge formation (control sample) lies in a range of . But for the other 3 samples, i.e. the ones showing signatures of bulge formation goes up to nearly . This is expected because bulge formation occurs in galaxies having a bar and high mass galaxies are more likely to have bars than the lower mass galaxies (Erwin (2018)).
We have defined axial ratio as the ratio of disk height to disk radius. For the galaxies having signature of bulge formation, we expect the axial ratio to be higher than that of the control sample as the disk height is higher for them. For our lognorm fit, we obtained the mean value of axial ratio of strong both-sided galaxies is higher among all samples (4b).
For all our further analysis described in section 7.4 and 7.5, we have used only the galaxies having mass, disk radius and disk height less than , 14kpc and 5kpc respectively. By using these limits, we were able to remove the relatively higher values in the data so that we could study the remaining data in a more detailed manner. These limits do not affect the correctness of the analysis because this includes nearly 98.8% of the whole galaxies. Category-wise, we have more than 99% of control sample and one-sided sample , nearly 98.3% of both-sided and 97% of strong both-sided inside this limit.
7.2 BP/X fraction
There have been a number of attempts to estimate the BP/X bulge fraction for different-size samples of edge-on galaxies. The earliest study was by Jarvis (1986), who estimated BP/X frequency of 1.2% in his relatively small sample of 41 disc galaxies. In a sample of 117 visually edge-on galaxies with diameter , Shaw (1987) estimated that approximately 20% contain a BP/X bulge. He noted this as a conservative lower limit, as additional BP/X bulge galaxies were identified that did not appear in the primary masking. Using a larger sample of 555 galaxies with axial ratio , De Souza & Dos Anjos (1987) estimated the BP/X bulge fraction to be 13%.
Lütticke et al. (2000) presented the first detailed statistics of BP/X bulges using a large sample of edge-on galaxies from Third Reference Catalogue of Bright Galaxies (RC3: de Vaucouleurs et al., 1991). They found BP/X bulges in their sample of 734 classifiable bulges. Yoshino & Yamauchi (2015) decomposed about 1700 SDSS DR7 disc galaxies, and identified BP/X structure. In their clean sample of 1329 galaxies in band, they reported BP/X structures. Since Yoshino & Yamauchi (2015) used SDSS sample, we cross-matched our sample with the their catalog of 1312 edge-on galaxies in the -band, and identified 439 galaxies in common. The primary reason for the discrepancy in sample sizes lies in the differing selection criteria.Yoshino & Yamauchi (2015) selected galaxies with a minor-to-major isophotal axis ratio in the -band less than 0.25, whereas our sample was selected based on volunteer votes. When we compare the values from their model fitting for both the common and complete samples, we find that the common sample predominantly consists of galaxies with , while complete sample show median , indicating that a large fraction of their sample comprises moderately inclined systems. Nonetheless, out of these common galaxies they identified 123 total BP/X structures (7 strong, 40 standard, and 76 weak in their definitions) and we identified 150 total BP/X structures (65 strong both-sided and 85 both-sided). Recently, a similar fraction of BP/X bulges is reported in -band images. Marchuk et al. (2022) trained Artificial Neural Network (ANN) to automatically identify BP/X structures in large sample of 13048 edge-on galaxies obtained from DESI Legacy Imaging Survey (Dey et al., 2019). They found about 1925 () galaxies with BP/X structures.
As described previously in Section 4, we started with 6653 edge-on galaxies and classified them into three classes of BP/X bulges: Both-sided, Strong both-sided, and One-sided. We identify 1675 () both-sided bulges, and 1108 () one-sided bulges. Among both-sided bulges 545 () are strong both-sided with prominent X-structure, and remaining 1130 () show faint X-structure. When combined together, both-sided and one-sided make (2783 galaxies) of the total sample. Though our base sample is nearly half of that studied in Marchuk et al. (2022), we have visually verified the presence of BP/X structure in each galaxy by manually varying intensity of the residual images to avoid missing any faint structure. Therefore, our analysis presents largest and complete sample of BP/X structures till date with detailed classification.
7.3 Relationship of BP/X bulge with the stellar mass
The fraction of barred galaxies shows a strong correlation with the stellar mass of a galaxy (Nair & Abraham, 2010; Melvin et al., 2014; Cervantes Sodi, 2017). Generally, massive galaxies tend to have bar more likely. Analogously, by considering a sample of 84 barred galaxies, Erwin & Debattista (2017) showed that the BP/X bulges are more common in galaxies with higher stellar masses. They found about 80% of barred galaxies with stellar masses have BP/X bulges. They concluded that the frequency of BP/X bulge in barred galaxies is nearly 50% in the stellar mass range . The similar stellar mass-dependence of BP/X bulges has been reported in other studies (Marchuk et al., 2022), and at higher redshift as well (Kruk et al., 2019).
In Fig. 4(a), we have plotted the stellar mass distribution of both-sided (gold), strong both-sided (green), one-sided (black), and control sample (cyan) in this study. We found that the distribution of control galaxies peaks at lower stellar mass than the galaxies with BP/X bulges. The median of the control sample lies at , whereas the sample with both-sided, strong both-sided, and one-sided BP/X bulges have median at , respectively. Our results are consistent with the previous findings of Erwin & Debattista (2017); Kruk et al. (2019); Marchuk et al. (2022), suggesting massive galaxies most likely host the BP/X bulges.
Further, we measure the BP/X fraction in our sample of edge-on galaxies. Fig. 5 shows the fraction of both-sided (green), strong both-sided (orange), one-sided (blue) bulges with the stellar mass of galaxies. The vertical bars in respective curve represent statistical uncertainty in measurement defined as , where being fraction of a bulge class at any stellar mass bin and being total number of galaxies in that bin. Similar to the previous studies, we notice increasing BP/X bulge fraction with increasing stellar mass of galaxies. In the stellar mass range shown here, the fraction of one-sided BP/X bulges is always higher than the strong both-sided bulges at any given stellar mass. Additionally, the fraction of both-sided BP/X bulges, which also includes strong both-sided bulges, is always the highest for any stellar mass. In the shown stellar mass range, we can see an increase in strong both-sided, one-sided and both-sided bulge fraction, respectively, from 0, 0, and 12 to 18, 24, and 38 percent.
7.4 Dependence of BP/X bulge on stellar surface density
To investigate the dependence of BP/X structures formation on the disk stellar surface density, we measure the average stellar surface density of galaxies in our sample. We consider that the total stellar mass of a galaxy lies within its Petrosian radius and define average stellar surface density as . Note that we consider Petrosian radius for measuring average stellar density instead of scale radius because measurement of scale radius is significantly affected by the secular evolution of inner disk and presence of dust as mentioned earlier in Section 3.3 (see Appendix B for similar analysis using scale radius). In Fig. 6, we compare the fraction of both-sided, one-sided, and strong both-sided BP/X structures as a function of average stellar surface density of galaxies, respectively, in green, red, and purple colors.
Fig. 6 clearly shows that the fraction of any type of BP/X structure increases with the increasing average surface density of galaxies. The low surface density disks () show nearly constant fraction of both-sided and strong both-sided BP/X bulges. Below this surface density limit, the fraction of strong both-sided structures is less than , while both-sided makes about , and the contribution of one-side structures rises from zero to . For high surface density disks (), the fraction of BP/X structure hosting disks rises by about in all three groups. This analysis shows that the high surface density disks are more likely to host BP/X structures as compare to the low surface density disks.
The increasing fraction of BP/X structures with increasing average surface density of galaxy provides additional constraint on the BP/X bulge formation. Though the fraction of BP/X structures strongly depends on the stellar mass of galaxies, this study suggests the domination of BP/X structures in high surface density disks. It is consistent with the bar formation criterion that the high surface density disk have low Toomre-Q parameter, which makes disk unstable against dynamical instability. An unstable dynamically cold disk can form bar easily, which can eventually thicken to form a BP/X structure (Combes & Sanders, 1981; Raha et al., 1991; Athanassoula, 2005).
7.5 Effect of BP/X structure on galaxy evolution
7.5.1 The mass-radius relation
The mass-radius relation of galaxies is a fundamental scaling law that links a galaxy’s stellar mass to its physical size. This relation shows that more massive galaxies tends to be larger in size. However, the slope and scatter of the relation depend on galaxy type. Early-type galaxies (ellipticals and lenticulars) typically exhibit a steeper mass–radius relation, indicating more compact structures at fixed mass compared to late-type (disk-dominated) galaxies (Shen et al., 2003; van der Wel et al., 2014). At a given stellar mass, early-type galaxies are smaller than late-type galaxies, reflecting differing formation mechanisms - such as gas-poor mergers for ellipticals and inside-out star formation for disks (Naab et al., 2009; Nelson et al., 2016). Moreover, high-redshift galaxies () are observed to be significantly more compact at fixed stellar mass than their local counterparts, suggesting substantial size growth over cosmic time (Trujillo et al., 2006; van Dokkum et al., 2010). Understanding the mass–radius relation is therefore critical for constraining galaxy assembly histories and testing models of galaxy evolution.
To investigate the role of galaxies hosting BP/X structures—and consequently the role of barred galaxies—in the mass–size relation of galaxies, we plot the Petrosian radius as a function of stellar mass for control sample and galaxies having strong both-sided BP/X bulges in the top panel of Fig. 7. We did not include both-sided and one-sided sample in this analysis to avoid any bias due to presence of heavy dust lanes, which may lead to their wrong classification in some cases. Also, we use Petrosian radius instead of scale radius to minimize the effect of dust obscuration and bar evolution on scale radius size. The control and strong both-sided samples are shown using blue, and purple color open circles, respectively. We have also plotted density contour for both the samples to show the spread of data using their respective colors. To show the size evolution with mass we have plot the median sizes in different mass bins along with 16th and 84th percentile ranges using large filled circles. The number of galaxies in each bin are listed at the bottom of the bottom panel figure. The straight lines are fit to median data points. To avoid the effect of redshift evolution, we have considered only those galaxies which lie within of the median redshift of our sample, i.e., to 0.09. We find that galaxies in the strong both-sided sample are typically smaller in size as compare to galaxies in control. Additionally, the disks with strong both-sided sample show shallower relation than control sample. This shallow trend contribute to large scatter at high-mass end of mass-size relation. However, the scatter in the sample is very high compare to the difference between to samples.
7.5.2 The mass-density relation
In analogous to mass-radius relations, the mass-density relations shows that more massive galaxies tend to have higher surface density. The early-type galaxies are more compact and centrally concentrated than the late-type galaxies. This correlation reflects the underlying processes governing star formation efficiency, feedback, and mass assembly. For instance, compact massive galaxies at high redshift exhibit high stellar surface densities, suggesting rapid, early star formation followed by quenching and morphological transformation (Barro et al., 2013).
We study the effect of BP/X hosting galaxies on the mass-density relation of spirals by fitting a straight line in log-log scale similar to the mass-size relation and plot in the bottom panel of Fig. 7. Again, blue color shows control sample and purple represents strong both-sided bulges. The mass-density relations are significantly stronger than the mass-size relations and shows higher degree of correlation. The galaxies having strong both-sided BP/X structures are typically compact towards high-mass end of the relation. The mass-density relation does not change as significantly as mass-size relation of galaxies.
8 Summary
In this paper, we performed a comprehensive study of BP/X structures in edge-on galaxies obtained from SDSS DR8. We compiled an initial sample of 6684 galaxies with more than 0.9 probability of being edge-on based on more than 10 votes in GalaxyZoo survey. We used GALFIT for 2D image decomposition to measure the scale radii and scale heights of galaxies. Further we used the residual images from 2D decomposition to manually identify BP/X structures in each image and categorized the sample in four classes depending on the signature of BP/X structure(1) strong both-sided: galaxies with clear BP/X structures, (2) both-sided: galaxies with faint or clear BP/X structures, (3) one-sided: galaxies with excess flux on one side on the galactic plan, and (4) control: galaxies without BP/X structures. Additionally, We calculated the mass, distance, and physical size of galaxies. Followings are key summarizing points of our study.
We visually classified 6653 edge-on galaxies into 1675 () both-sided bulges, and 1108 () one-sided bulges. Among both-sided bulges 545 () are strong both-sided with prominent X-structure, and remaining 1130 () show faint X-structure. When combined together, both-sided and one-sided make (2783 galaxies) of the total sample. Our sample provides the largest catalog of BP/X bulges to date, which includes a category of one-sided BP/X bulges. The catalog of one-sided structures is one of its kind sample, which likely contains samples of ongoing buckling events in edge-on galaxies.
The fraction of one-sided BP/X bulges is always higher than the strong both-sided bulges at any given stellar mass. The fraction of both-sided BP/X bulges is always the highest for any stellar mass. At , the fraction of strong both-sided, one-sided and both-sided bulge is 18, 24, and 38 percent, respectively.
We report a strong dependence of BP/X fraction on average stellar surface density of galaxy. The fraction of BP/X hosting galaxies increases with the surface density, reflecting bar formation in high surface density disk and subsequent BP/X structure growth.
We use our catalog to investigate their role on stellar mass-size relation and stellar mass-stellar surface density relation. We found that galaxies with strong both-sided BP/X structures contribute to the scatter in these relations at high-mass end. However, the scaling relations of control and strong both-sided BP/X show significant overlap in their scatter.
This study provides a classification of BP/X bulges in a large sample of edge-on galaxies, along with their statistical properties and evolutionary trends. We demonstrate the effect of bar formation and evolution on the mass–size and mass–density scaling relations. The analysis uses single-component fitting for size measurements and empirical relations for dust correction and mass estimation, which introduce significant scatter in the scaling relations. Future studies employing detailed 2D decomposition for scale size measurements and SED fitting for stellar mass, star formation rate, dust content, and star formation history will be instrumental in the era of large-scale surveys, e.g., Euclid, LSST, 4MOST, and Roman. Additionally, our sample can be used to train AI models for studying large datasets.
Acknowledgements.
We thank anonymous reviewer for providing constructive feedback and suggestion, which helped improve this paper. AK acknowledges support from ALMA fund with code 31220021 and from ANID BASAL project FB210003. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, The Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University. This publication uses data generated via the zooniverse.org platform, development of which is funded by generous support, including a Global Impact Award from Google, and by a grant from the Alfred P. Sloan Foundation. We make use of Matplotlib (Hunter, 2007), Numpy (Harris et al., 2020), Pandas (Wes McKinney, 2010; pandas development team, 2020), Scipy (Virtanen et al., 2020b).References
- Aguerri et al. (2009) Aguerri, J., Méndez-Abreu, J., & Corsini, E. 2009, Astronomy & Astrophysics, 495, 491
- Aihara et al. (2011) Aihara, H., Prieto, C. A., An, D., et al. 2011, The Astrophysical Journal Supplement Series, 193, 29
- Alonso et al. (2013) Alonso, M. S., Coldwell, G., & Lambas, D. G. 2013, A&A, 549, A141
- Athanassoula (2003) Athanassoula, E. 2003, Monthly Notices of the Royal Astronomical Society, 341, 1179
- Athanassoula (2005) Athanassoula, E. 2005, Monthly Notices of the Royal Astronomical Society, 358, 1477
- Athanassoula & Martinez-Valpuesta (2008) Athanassoula, E. & Martinez-Valpuesta, I. 2008, in Chaos in Astronomy (Springer), 77–83
- Baba & Kawata (2020) Baba, J. & Kawata, D. 2020, Monthly Notices of the Royal Astronomical Society, 492, 4500
- Barro et al. (2013) Barro, G., Faber, S. M., Pérez-González, P. G., et al. 2013, ApJ, 765, 104
- Bell et al. (2003) Bell, E. F., McIntosh, D. H., Katz, N., & Weinberg, M. D. 2003, The Astrophysical Journal Supplement Series, 149, 289
- Berentzen et al. (2007) Berentzen, I., Shlosman, I., Martinez-Valpuesta, I., & Heller, C. H. 2007, ApJ, 666, 189
- Bettoni & Galletta (1994) Bettoni, D. & Galletta, G. 1994, A&A, 281, 1
- Blanton et al. (2003) Blanton, M. R., Hogg, D. W., Bahcall, N. A., et al. 2003, The Astrophysical Journal, 592, 819
- Boone et al. (2007) Boone, F., Baker, A. J., Schinnerer, E., et al. 2007, A&A, 471, 113
- Bradley & Developers (2023) Bradley, L. & Developers, P. 2023, astropy/photutils: 1.6.0
- Bureau & Athanassoula (2005) Bureau, M. & Athanassoula, E. 2005, The Astrophysical Journal, 626, 159
- Cervantes Sodi (2017) Cervantes Sodi, B. 2017, ApJ, 835, 80
- Ciambur & Graham (2016) Ciambur, B. C. & Graham, A. W. 2016, Monthly Notices of the Royal Astronomical Society, 459, 1276
- Cole et al. (2014) Cole, D. R., Debattista, V. P., Erwin, P., Earp, S. W., & Roškar, R. 2014, Monthly Notices of the Royal Astronomical Society, 445, 3352
- Collier (2020) Collier, A. 2020, MNRAS, 492, 2241
- Combes et al. (1990) Combes, F., Debbasch, F., Friedli, D., & Pfenniger, D. 1990, Astronomy and Astrophysics (ISSN 0004-6361), vol. 233, no. 1, July 1990, p. 82-95. Research supported by the Universite de Geneve and SNSF., 233, 82
- Combes & Sanders (1981) Combes, F. & Sanders, R. 1981, Astronomy and Astrophysics, vol. 96, no. 1-2, Mar. 1981, p. 164-173., 96, 164
- de Sá-Freitas et al. (2023) de Sá-Freitas, C., Fragkoudi, F., Gadotti, D. A., et al. 2023, A&A, 671, A8
- De Souza & Dos Anjos (1987) De Souza, R. & Dos Anjos, S. 1987, Astronomy and Astrophysics Supplement Series, 70, 465
- de Vaucouleurs et al. (1991) de Vaucouleurs, G., de Vaucouleurs, A., Corwin, Herold G., J., et al. 1991, Third Reference Catalogue of Bright Galaxies
- Debattista (2018) Debattista, V. 2018, in The Galactic Bulge at the Crossroads, 10
- Debattista et al. (2020) Debattista, V. P., Liddicott, D. J., Khachaturyants, T., & Beraldo e Silva, L. 2020, MNRAS, 498, 3334
- Debattista et al. (2006) Debattista, V. P., Mayer, L., Carollo, C. M., et al. 2006, ApJ, 645, 209
- Dey et al. (2019) Dey, A., Schlegel, D. J., Lang, D., et al. 2019, AJ, 157, 168
- Di Matteo et al. (2019) Di Matteo, P., Haywood, M., Lehnert, M., et al. 2019, Astronomy & Astrophysics, 632, A4
- D´ıaz-Garc´ıa et al. (2020) Díaz-García, S., Moyano, F. D., Comerón, S., et al. 2020, A&A, 644, A38
- D´ıaz-Garc´ıa et al. (2016) Díaz-García, S., Salo, H., Laurikainen, E., & Herrera-Endoqui, M. 2016, Astronomy & Astrophysics, 587, A160
- Donohoe-Keyes et al. (2019) Donohoe-Keyes, C., Martig, M., James, P., & Kraljic, K. 2019, Monthly Notices of the Royal Astronomical Society, 489, 4992
- Du & McGaugh (2020) Du, W. & McGaugh, S. S. 2020, The Astronomical Journal, 160, 122
- Erwin (2018) Erwin, P. 2018, MNRAS, 474, 5372
- Erwin & Debattista (2016) Erwin, P. & Debattista, V. P. 2016, The Astrophysical Journal Letters, 825, L30
- Erwin & Debattista (2017) Erwin, P. & Debattista, V. P. 2017, Monthly Notices of the Royal Astronomical Society, 468, 2058
- Faúndez-Abans & de Oliveira-Abans (1998) Faúndez-Abans, M. & de Oliveira-Abans, M. 1998, Astronomy and Astrophysics Supplement Series, 128, 289
- Fragkoudi et al. (2017) Fragkoudi, F., Di Matteo, P., Haywood, M., et al. 2017, Astronomy & Astrophysics, 606, A47
- Fraser-McKelvie et al. (2020) Fraser-McKelvie, A., Aragón-Salamanca, A., Merrifield, M., et al. 2020, MNRAS, 495, 4158
- Friedli & Benz (1993) Friedli, D. & Benz, W. 1993, Astronomy and Astrophysics (ISSN 0004-6361), vol. 268, no. 1, p. 65-85., 268, 65
- Friedli & Benz (1995) Friedli, D. & Benz, W. 1995, Astronomy and Astrophysics, v. 301, p. 649, 301, 649
- Friedli & Pfenniger (1990) Friedli, D. & Pfenniger, D. 1990, in European Southern Observatory Conference and Workshop Proceedings, Vol. 35, European Southern Observatory Conference and Workshop Proceedings, ed. B. J. Jarvis & D. M. Terndrup, 265
- Galloway et al. (2015) Galloway, M. A., Willett, K. W., Fortson, L. F., et al. 2015, MNRAS, 448, 3442
- Ghosh et al. (2024) Ghosh, S., Fragkoudi, F., Di Matteo, P., & Saha, K. 2024, A&A, 683, A196
- Ginsburg et al. (2019) Ginsburg, A., Sipőcz, B. M., Brasseur, C. E., et al. 2019, AJ, 157, 98
- Gonzalez et al. (2017) Gonzalez, O. A., Debattista, V. P., Ness, M., Erwin, P., & Gadotti, D. A. 2017, Monthly Notices of the Royal Astronomical Society: Letters, 466, L93
- Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357–362
- Heller & Shlosman (1994) Heller, C. H. & Shlosman, I. 1994, Astrophysical Journal, Part 1 (ISSN 0004-637X), vol. 424, no. 1, p. 84-105, 424, 84
- Hunter (2007) Hunter, J. D. 2007, Computing in Science Engineering, 9, 90
- Jarvis (1986) Jarvis, B. J. 1986, The Astronomical Journal, 91, 65
- Kalnajs (1972) Kalnajs, A. J. 1972, ApJ, 175, 63
- Kataria & Vivek (2024) Kataria, S. K. & Vivek, M. 2024, MNRAS, 527, 3366
- Kautsch et al. (2006) Kautsch, S., Grebel, E., Barazza, F., & Gallagher, J. 2006, Astronomy & Astrophysics, 445, 765
- Khoperskov et al. (2019) Khoperskov, S., Di Matteo, P., Gerhard, O., et al. 2019, A&A, 622, L6
- Khoperskov et al. (2018) Khoperskov, S., Haywood, M., Di Matteo, P., Lehnert, M., & Combes, F. 2018, Astronomy & Astrophysics, 609, A60
- Kim et al. (2011) Kim, S. S., Saitoh, T. R., Jeon, M., et al. 2011, The Astrophysical Journal Letters, 735, L11
- Kruk et al. (2019) Kruk, S. J., Erwin, P., Debattista, V. P., & Lintott, C. 2019, MNRAS, 490, 4721
- Kuijken & Merrifield (1995) Kuijken, K. & Merrifield, M. R. 1995, arXiv preprint astro-ph/9501114
- Kumar (2023) Kumar, A. 2023, arXiv e-prints, arXiv:2306.11045
- Kumar et al. (2021) Kumar, A., Das, M., & Kataria, S. K. 2021, MNRAS, 506, 98
- Kumar et al. (2022) Kumar, A., Das, M., & Kataria, S. K. 2022, Monthly Notices of the Royal Astronomical Society, 509, 1262
- Kumar & Kataria (2022) Kumar, A. & Kataria, S. K. 2022, Monthly Notices of the Royal Astronomical Society, 514, 2497
- Lang et al. (2014) Lang, M., Holley-Bockelmann, K., & Sinha, M. 2014, ApJ, 790, L33
- Laurikainen & Salo (2016) Laurikainen, E. & Salo, H. 2016, in Galactic Bulges (Springer), 77–106
- Li et al. (2017) Li, Z.-Y., Ho, L. C., & Barth, A. J. 2017, ApJ, 845, 87
- Lintott et al. (2010) Lintott, C., Schawinski, K., Bamford, S., et al. 2010, Monthly Notices of the Royal Astronomical Society, 410, 166
- Lintott et al. (2008) Lintott, C. J., Schawinski, K., Slosar, A., et al. 2008, Monthly Notices of the Royal Astronomical Society, 389, 1179
- Łokas (2018) Łokas, E. L. 2018, ApJ, 857, 6
- Łokas (2019) Łokas, E. L. 2019, Astronomy & Astrophysics, 629, A52
- Lütticke et al. (2000) Lütticke, R., Dettmar, R.-J., & Pohlen, M. 2000, Astronomy and Astrophysics Supplement Series, 145, 405
- Lynden-Bell (1979) Lynden-Bell, D. 1979, Monthly Notices of the Royal Astronomical Society, 187, 101
- Marchuk et al. (2022) Marchuk, A. A., Smirnov, A. A., Sotnikova, N. Y., et al. 2022, MNRAS, 512, 1371
- Marchuk et al. (2022) Marchuk, A. A., Smirnov, A. A., Sotnikova, N. Y., et al. 2022, Monthly Notices of the Royal Astronomical Society, 512, 1371
- Martin & Friedli (1997) Martin, P. & Friedli, D. 1997, Astronomy and Astrophysics, v. 326, p. 449-464, 326, 449
- Martinez-Valpuesta et al. (2006) Martinez-Valpuesta, I., Shlosman, I., & Heller, C. 2006, ApJ, 637, 214
- Masters et al. (2011) Masters, K. L., Nichol, R. C., Hoyle, B., et al. 2011, Monthly Notices of the Royal Astronomical Society, 411, 2026
- Melvin et al. (2014) Melvin, T., Masters, K., Lintott, C., et al. 2014, MNRAS, 438, 2882
- Merritt & Sellwood (1994) Merritt, D. & Sellwood, J. 1994, Astrophysical Journal, Part 1 (ISSN 0004-637X), vol. 425, no. 2, p. 551-567, 425, 551
- Minchev et al. (2011) Minchev, I., Famaey, B., Combes, F., et al. 2011, A&A, 527, A147
- Miwa & Noguchi (1998) Miwa, T. & Noguchi, M. 1998, ApJ, 499, 149
- Naab et al. (2009) Naab, T., Johansson, P. H., & Ostriker, J. P. 2009, ApJ, 699, L178
- Nair & Abraham (2010) Nair, P. B. & Abraham, R. G. 2010, ApJ, 714, L260
- Nelson et al. (2016) Nelson, E. J., van Dokkum, P. G., Förster Schreiber, N. M., et al. 2016, ApJ, 828, 27
- Ness & Lang (2016) Ness, M. & Lang, D. 2016, AJ, 152, 14
- Ostriker & Peebles (1973) Ostriker, J. P. & Peebles, P. J. E. 1973, ApJ, 186, 467
- pandas development team (2020) pandas development team, T. 2020, pandas-dev/pandas: Pandas
- Peng et al. (2010) Peng, C. Y., Ho, L. C., Impey, C. D., & Rix, H.-W. 2010, AJ, 139, 2097
- Pfenniger & Friedli (1991) Pfenniger, D. & Friedli, D. 1991, Astronomy and Astrophysics (ISSN 0004-6361), vol. 252, no. 1, Dec. 1991, p. 75-93. Research supported by Observatoire de Geneve and SNSF., 252, 75
- Polyachenko & Polyachenko (2003) Polyachenko, V. & Polyachenko, E. 2003, Astronomy Letters, 29, 447
- Quillen (2002) Quillen, A. C. 2002, The Astronomical Journal, 124, 722
- Quillen et al. (2014) Quillen, A. C., Minchev, I., Sharma, S., Qin, Y.-J., & Di Matteo, P. 2014, Monthly Notices of the Royal Astronomical Society, 437, 1284
- Raha et al. (1991) Raha, N., Sellwood, J., James, R., & Kahn, F. 1991, nature, 352, 411
- Saha et al. (2013) Saha, K., Pfenniger, D., & Taam, R. E. 2013, The Astrophysical Journal, 764, 123
- Savchenko et al. (2017) Savchenko, S., Sotnikova, N. Y., Mosenkov, A., Reshetnikov, V., & Bizyaev, D. 2017, Monthly Notices of the Royal Astronomical Society, 471, 3261
- Savchenko et al. (2023) Savchenko, S. S., Poliakov, D. M., Mosenkov, A. V., et al. 2023, MNRAS, 524, 4729
- Sellwood & Gerhard (2020) Sellwood, J. & Gerhard, O. 2020, Monthly Notices of the Royal Astronomical Society, 495, 3175
- Sellwood (2014) Sellwood, J. A. 2014, Reviews of Modern Physics, 86, 1
- Seo et al. (2019) Seo, W.-Y., Kim, W.-T., Kwak, S., et al. 2019, The Astrophysical Journal, 872, 5
- Shao et al. (2007) Shao, Z., Xiao, Q., Shen, S., et al. 2007, The Astrophysical Journal, 659, 1159
- Shaw (1987) Shaw, M. A. 1987, Monthly Notices of the Royal Astronomical Society, 229, 691
- Shen et al. (2003) Shen, S., Mo, H. J., White, S. D. M., et al. 2003, MNRAS, 343, 978
- Shlosman et al. (1989) Shlosman, I., Frank, J., & Begelman, M. C. 1989, Nature, 338, 45
- Silva-Lima et al. (2022) Silva-Lima, L. A., Martins, L. P., Coelho, P. R. T., & Gadotti, D. A. 2022, A&A, 661, A105
- Simard et al. (2011) Simard, L., Mendel, J. T., Patton, D. R., Ellison, S. L., & McConnachie, A. W. 2011, The Astrophysical Journal Supplement Series, 196, 11
- Smirnov & Sotnikova (2018) Smirnov, A. A. & Sotnikova, N. Y. 2018, Monthly Notices of the Royal Astronomical Society, 481, 4058
- Smirnov & Sotnikova (2019) Smirnov, A. A. & Sotnikova, N. Y. 2019, MNRAS, 485, 1900
- Sormani et al. (2022) Sormani, M. C., Gerhard, O., Portail, M., Vasiliev, E., & Clarke, J. 2022, Monthly Notices of the Royal Astronomical Society: Letters, 514, L1
- Spinoso et al. (2016) Spinoso, D., Bonoli, S., Dotti, M., et al. 2016, Monthly Notices of the Royal Astronomical Society, 465, 3729
- Stoughton et al. (2002) Stoughton, C., Lupton, R. H., Bernardi, M., et al. 2002, The Astronomical Journal, 123, 485
- Stoughton et al. (2002) Stoughton, C., Lupton, R. H., Bernardi, M., et al. 2002, AJ, 123, 485
- Toomre (1981) Toomre, A. 1981, in Structure and Evolution of Normal Galaxies, ed. S. M. Fall & D. Lynden-Bell, 111–136
- Trujillo et al. (2006) Trujillo, I., Förster Schreiber, N. M., Rudnick, G., et al. 2006, ApJ, 650, 18
- Van der Kruit & Searle (1981) Van der Kruit, P. & Searle, L. 1981, Astronomy and Astrophysics, vol. 95, no. 1, Feb. 1981, p. 105-115., 95, 105
- van der Kruit & Freeman (2011) van der Kruit, P. C. & Freeman, K. C. 2011, ARA&A, 49, 301
- van der Wel et al. (2014) van der Wel, A., Franx, M., van Dokkum, P. G., et al. 2014, ApJ, 788, 28
- van Dokkum et al. (2010) van Dokkum, P. G., Whitaker, K. E., Brammer, G., et al. 2010, ApJ, 709, 1018
- Veilleux et al. (1999) Veilleux, S., Bland-Hawthorn, J., & Cecil, G. 1999, AJ, 118, 2108
- Virtanen et al. (2020a) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020a, Nature Methods, 17, 261
- Virtanen et al. (2020b) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020b, Nature Methods, 17, 261
- Wes McKinney (2010) Wes McKinney. 2010, in Proceedings of the 9th Python in Science Conference, ed. Stéfan van der Walt & Jarrod Millman, 56 – 61
- Wozniak (2007) Wozniak, H. 2007, Astronomy & Astrophysics, 465, L1
- Xiang et al. (2021) Xiang, K. M., Nataf, D. M., Athanassoula, E., et al. 2021, ApJ, 909, 125
- York et al. (2000) York, D. G., Adelman, J., Anderson, Jr., J. E., et al. 2000, AJ, 120, 1579
- Yoshino & Yamauchi (2015) Yoshino, A. & Yamauchi, C. 2015, Monthly Notices of the Royal Astronomical Society, 446, 3749
- Yuan et al. (2013) Yuan, H.-B., Liu, X.-W., & Xiang, M.-S. 2013, Monthly Notices of the Royal Astronomical Society, 430, 2188
- Zozulia et al. (2024a) Zozulia, V. D., Smirnov, A. A., & Sotnikova, N. Y. 2024a, Monthly Notices of the Royal Astronomical Society, 529, 4405
- Zozulia et al. (2024b) Zozulia, V. D., Smirnov, A. A., Sotnikova, N. Y., & Marchuk, A. A. 2024b, Astronomy & Astrophysics, 692, A145
- Zozulia et al. (2025) Zozulia, V. D., Sotnikova, N. Y., & Smirnov, A. A. 2025, Astronomy & Astrophysics, 698, L26
Appendix A More examples of sample galaxies
In order to provide a clear overview of the different categories of galaxies with BP/X bulges defined in this study, we present mosaic images for each category. Fig. 8 shows a mosaic of strong, both-sided BP/X bulges; Fig. 9 displays a mosaic of weak, both-sided BP/X bulges; and Fig. 10 exhibits a mosaic of one-sided BP/X bulges. Each panel in these images shows the residuals obtained by subtracting the model images from the observed ones. The color scales in each panel are adjusted to enhance the visibility of structural features.
Appendix B Effect of measuring surface density using scale radius
In Section 7.4, we discussed the dependence of the BP/X bulge-hosting galaxy fraction on the average stellar surface density, measured by assuming all stellar mass lies within the Petrosian radius. Here, we examine the effect of using the disk scale radius instead for calculating the average stellar density. Specifically, we compute the average stellar surface density under the assumption that all stellar mass is contained within five scale radii, and define it as . Using a larger or smaller number of scale radii for this calculation does not affect the analysis qualitatively, as it merely scales the density by a constant factor. In Fig. 11, we present the BP/X fraction as a function of for three bulge categories. For comparison with Fig. 6, we limit the average density range in the range of that shown in Fig. 6, scaled by a factor of , which corresponds to the average value of . The results are highly consistent with those presented in Section 7.4, showing that the fraction of BP/X bulges increases with increasing stellar surface density. There are very small differences at low and high surface density ends likely due to small sample size.
Similar to the mass–density relation discussed in Section 7.5, which assumes total mass within the Petrosian radius, we investigate the effect of using the scale radius in Fig. 12. Here, we show the mass–size relation in the top panel and the mass–density relation in the bottom panel, using the scale radius instead of the Petrosian radius. In contrast to the trends in Fig. 7, both the control and strong both-sided BP/X bulge samples now show similar trends. This is likely due to the effects of bar thickening and dust distribution in galaxies, which influence scale radius measurements.

### v2

@@
 Residual images are visually inspected to classify BP/X features into four categories: strong both-sided, both-sided, one-sided, and control (no BP/X).
-We also estimated stellar mass, distance, and physical size for each galaxy.
+We also estimate stellar mass, distance, and physical size for each galaxy.
 Results.
-Out of 6653 classified galaxies, we identified 1675 (25%) with both-sided BP/X features—545 (8%) strong and 1130 (17%) faint—as well as 1108 (17%) one-sided structures, making up a total of 2783 BP/X-hosting galaxies (42%).
+Out of 6653 classified galaxies, we identified 1673 (25%) with both-sided BP/X features—504 () strong and 1169 () weak—as well as 1112 () one-sided structures, making up a total of 2785 BP/X-hosting galaxies ().
 One-sided structures, likely signatures of ongoing buckling, are more frequent than strong both-sided bulges across all stellar masses.
@@
 Key Words.:
-Galaxies: formation – Galaxies: evolution1 Introduction
-Spiral galaxies with elongated bar-like structures in their center are commonly known as barred galaxies and comprise nearly 30-70% of all spiral galaxies (Aguerri et al., 2009; Masters et al., 2011; D´ıaz-Garc´ıa et al., 2016).
+Galaxies: formation – Galaxies: evolution – Galaxies: bulges – Galaxies: structure – Galaxies: photometry – Galaxies: statistics1 Introduction
+Spiral galaxies with elongated bar-like structures in their center are commonly known as barred galaxies and comprise nearly 30-70% of all spiral galaxies (Aguerri et al., 2009; Masters et al., 2011; Díaz-García et al., 2016).
 Bars play a major role in the evolution of spiral galaxies (Marchuk et al., 2022).
 They are thought to funnel a large amount of gas into the nuclear region (Shlosman et al., 1989; Boone et al., 2007).
-This gas helps in the formation of new stars in the central and bar regions (D´ıaz-Garc´ıa et al., 2020; Fraser-McKelvie et al., 2020).
+This gas helps in the formation of new stars in the central and bar regions (Díaz-García et al., 2020; Fraser-McKelvie et al., 2020).
 The accretion of gas into the nuclear region also feeds the central supermassive black holes (SMBHs) that help the growth of SMBHs.
-Sometimes the excess gas accretion onto SMBH ignites the nuclear activity in the galaxy (Alonso et al., 2013; Galloway et al., 2015; Silva-Lima et al., 2022; Kataria & Vivek, 2024).
+Sometimes the excess gas accretion onto SMBH ignites the nuclear activity in the galaxy (Alonso et al., 2013; Galloway et al., 2015; Silva-Lima et al., 2022; Garland et al., 2024; Kataria & Vivek, 2024).
 During the growth phase of bars, nuclear stellar disc (NSD) is formed due to intense star formation in the central sub-kpc region (Friedli & Benz, 1993; Heller & Shlosman, 1994; Friedli & Benz, 1995; Martin & Friedli, 1997; Wozniak, 2007; Kim et al., 2011; Cole et al., 2014; Seo et al., 2019; Baba & Kawata, 2020; Sormani et al., 2022).
 The bar suppresses star formation through out the remainder of its extent (Martin & Friedli, 1997; Spinoso et al., 2016; Khoperskov et al., 2018; Donohoe-Keyes et al., 2019).
-Therefore, the NSD has younger stars than the age of the bar (Baba & Kawata, 2020) and the BPX bulge contains stars older than the NSD (Gonzalez et al., 2017).
+Therefore, the NSD has younger stars than the age of the bar (Baba & Kawata, 2020).
 Using their numerical simulations, Baba & Kawata (2020) showed that the oldest stellar population in the NSD can be used for dating bar formation time (also see de Sá-Freitas et al., 2023).
@@
 However, it is not easy to detect the presence of a BP/X bulge, as it requires the bulge to be oriented in a perpendicular direction with respect to our line of sight.
-There has been an attempt to create a catalogue of edge-on galaxies (Kautsch et al., 2006).
+There has been an attempt to create a catalog of edge-on galaxies (Kautsch et al., 2006).
 Additionally, there have been attempts to study the relationship between the development of BP/X bulges and the dark matter halo using numerical simulations (Kumar et al., 2022).
@@
 However, the current gas mass fraction does not appear to influence whether a BP/X bulge will form (Erwin & Debattista, 2017).
-Here, we aim to separate the edge-on galaxies having BP/X bulge from a large sample of edge-on galaxies obtained from the Sloan Digital Sky Survey111SDSS.
+Here, we aim to separate the edge-on galaxies having BP/X bulge from a large sample of edge-on galaxies obtained from the Sloan Digital Sky Survey.
 We classify galaxies having BP/X bulges and study different parameters like masses and sizes of galaxies with or without BP/X bulges.
-Then we compare their properties to find any speciality in the sample with BP/X bulges.
+Then we compare their properties to find any specialty in the sample with BP/X bulges.
 We verified some of the past similar analysis performed previously on relatively smaller sample size.
@@
 Followings are the two criteria we have imposed to select desired galaxies:
-1.
 Number of votes (nvote) ,
-2.
 Fraction of votes for edge-on galaxy (p_edge) ¿ 0.9
@@
 A fraction greater than 0.9 signifies that the most volunteers have identified the galaxy as edge-on.
-Then we obtain coordinates (RA, Dec) of the selected galaxies from the ”Galaxy” table of the Sloan Digital Sky Survey data release 8 (SDSS DR8)222SDSS DR8.
+Then we obtain coordinates (RA, Dec) of the selected galaxies from the ”Galaxy” table of the Sloan Digital Sky Survey data release 8 (SDSS DR8)111SDSS DR8.
 The ”Galaxy” table contains photometric data of all the objects classified as galaxies from the ”PhotoPrimary” table.
@@
 The SQL search query for this task is attached below.
-SELECT g.ra, g.dec, g.sky_r, g.petroRad_r, g.sky_r, g.expPhi_r FROM Galaxy g JOIN zooSpec z ON z.objid = g.objID WHERE z.nvote > 10 AND z.p_edge > 0.90
+SELECT
+g.ra, g.dec, g.sky_r, g.petroRad_r,
+g.sky_r, g.expPhi_r
+FROM
+Galaxy g
+JOIN zooSpec z ON z.objid = g.objID
+WHERE
+z.nvote > 10 AND
+z.p_edge > 0.90
 2.2 FITS image files
@@
 Our choice of -band images reflects our interest in the stellar light profile of galaxies because -band is less affected by dust extinction and has higher S/N ratio (York et al., 2000), which makes it easier to distinguish other structures from the oval (Yoshino & Yamauchi, 2015).
-For this purpose, we utilise the module 333Skyview of the python package Astroquery (Ginsburg et al., 2019).
+For this purpose, we utilize the SDSS module 222astroquery.SDSS of the Astroquery (Ginsburg et al., 2019) Python package.
 Astroquery provide a set of tools for accessing astronomical data from the webs of various surveys.
@@
 The pre-computed Petrosian radii of the galaxies are obtained using SDSS SQL search query along with the (RA, Dec) positions as discussed in Section 2.1.
-For all the galaxies, we used the default Skyview image dimensions, i.e., 300 pixels 300 pixels.
-Due to our default image dimensions, the pixel size varies among galaxies depending upon the corresponding petrosian radius.
-Out of 9068 galaxies, we can obtain images for 6684 galaxies because Skyview does not return images for some specific positions.
+For all the galaxies, we used the default SDSS pixel scale, i.e.
+arcsec/pixel.
+Out of 9068 galaxies, we can obtain images for 6684 galaxies because SDSS does not return images for some specific positions.
 Through a quick visual inspection we removed incomplete images (where only some part of the galaxy is visible) leaving a sample of 6653 galaxies for further analysis.
@@
 These parameters include object ID (objid), redshift (z), absolute magnitudes.
-For this purpose we make use of (RA, DEC) positions of selected galaxies in SDSS crossID tool 444SDSS crossID.
+For this purpose we make use of (RA, DEC) positions of selected galaxies in SDSS crossID tool 333SDSS crossID.
 This tool allows users to upload their own table which can be joined with different SDSS tables to extract other information using a similar SQL query as described in Section 2.1.
@@
 Using the median or Gaussian filter to smooth the image, and then subtracting these smoothed image from original image (and sometimes adding a factor of residual image back into the original image) to sharp the hidden BP/X structure has been used in the literature (Faúndez-Abans & de Oliveira-Abans, 1998).
-Therefore, we first started with using the median filter available in the Scipy Python package555Scipy (Virtanen et al., 2020a) to identify the BP/X hosting galaxies.
+Therefore, we first started with using the median filter available in the Scipy Python package444Scipy (Virtanen et al., 2020a) to identify the BP/X hosting galaxies.
 The median filter smooth any image according to input box (or window) size.
@@
 The subtraction of smooth intensity distribution of galaxy from its observed intensity distribution can also sharpen the hidden BP/X structure.
-Therefore, we construct smooth intensity distribution of galaxy by fitting iso-intensity ellipses using a Python package Photutils 666ellipse.photutils (Bradley & Developers (2023)).
+Therefore, we construct smooth intensity distribution of galaxy by fitting iso-intensity ellipses using a Python package Photutils 555ellipse.photutils (Bradley & Developers (2023)).
 It requires an initial guess for an ellipse geometry to start the fitting.
@@
 Our wrapping algorithm for automation takes care of the variable parameters, e.g., initial guesses, galaxy name, output file name, etc.
-This fitting analysis provides us with edge-disk parameters which are not available in the original SDSS catalogue.
+This fitting analysis provides us with edge-disk parameters which are not available in the original SDSS catalog.
 In Fig.
@@
 4 BP/X bulge classification
-After edge-disk fitting using galfit, we visually inspected residual images ( original image model image) of all 6653 galaxies to build a catalogue of galaxies with and without BP/X bulges for detailed study.
+After edge-disk fitting using galfit, we visually inspected residual images ( original image model image) of all 6653 galaxies to build a catalog of galaxies with and without BP/X bulges for detailed study.
 The sample without BP/X bulges is considered as control sample for comparison with the sample having BP/X bulges.
@@
 Following is our classification scheme for BP/X bulges and control sample.
-Both-sided:- This class of galaxies show the presence of x-shape or peanut-shape structure at their center in the residual images.
+Both-sided:- This class of galaxies show the presence of boxy, peanut, or x-shape structure at their center in the residual images.
 The structure can be clearly visible as in Fig.
@@
 2(d,e,f).
-We identify 1675 galaxies in this class, which comprise nearly of the total sample.
+We identify 1673 galaxies in this class, which comprise nearly of the total sample.
+On the basis of visibility, this class can be further divided into strong and weak both-sided structures.
+In our analysis, we refer to both-sided as the collection of both strong and weak both-sided structures, unless specified differently.
 Strong both-sided:- From the class of both-sided BP/X bulge galaxies, we separate galaxies showing a clear x-shape structures and classify them into strong both-sided sample.
@@
 2(a,b,c).
-There are 545 galaxies (nearly ) in this class.
+There are 504 galaxies (nearly ) in this class.
 One-sided:- The residual images of these galaxies shows excess intensity on one side of the galactic mid-plane.
@@
 Additionally, it is also possible that this class potentially contains galaxies in the initial buckling phase.
-There are 1108 galaxies in this class, which is nearly of the total sample.
+There are 1112 galaxies in this class, which is nearly of the total sample.
 Control:- These galaxies show no sign of BP/X bulge formation.
 Therefore, the model completely fits the original image, leaving no feature in the residual.
-There are a total of 3870 galaxies in this class.
+There are a total of 3868 galaxies in this class.
 Some of the galaxies in the control sample show presence of other significantly bright structure in the projection, e.g., in the bottom left side of Fig.
@@
 There are 146 galaxies which show nearby bright structures (likely neighbor galaxy).
-The remaining 3724 galaxies do not show any neighboring bright objects in the projection.
+The remaining 3722 galaxies do not show any neighboring bright objects in the projection.
 So, nearly of our sample belongs to this class.
 Note that some of the galaxies showing showing spheroidal structures (classical bulge or end-on projected bars) are also included in the control sample.
+For more examples of different categories of BP/X bulges discussed here, we refer the reader to Fig.
+8, 9, and 10 in Appendix A.
+We note that the weak both-sided and one-sided structures may contain some oblate classical bulge structures, which are difficult to distinguish due to spatial resolution limit of the SDSS.
+We caution the user to consider this caveat when interpreting results for these two classes of BP/X structures in this study.
 5 Derived quantities for galaxies
-Along with the parameters obtained from SDSS catalogue and measured using 2D image decomposition using galfit, we derive stellar masses, distances and physical sizes of the galaxies from existing information.
-The following subsections describes our methodology for measuring these derived properties of galaxies.
+Along with the parameters obtained from SDSS catalog and measured using 2D image decomposition using galfit, we derive stellar masses, distances and physical sizes of the galaxies from existing information.
+The following subsections describe our methodology for measuring these derived properties of galaxies.
 5.1 Stellar masses
@@
 We use the extinction corrected absolute magnitudes and colors for mass calculation.
-After calculating mass-to-light ratio using above described method, we use absolute magnitudes (Mr) to measure the luminosities of galaxies in solar unit by following expression.
-where, M⊙ is solar magnitude in .
+After calculating mass-to-light ratio using above described method, we use absolute -band magnitudes (Mr) to measure the luminosities of galaxies in solar unit by following expression.
+where, M⊙ is solar magnitude in -band.
 For our analysis, we obtained solar magnitude from Blanton et al.
@@
 From the galfit photometric decomposition, we obtain the disk scale height and disk scale radius in pixels.
-Since we have downloaded the FITS images of galaxies within 3 Petrosian radius aperture with 300 pixels 300 pixels size, the pixel size is simply Petrosian radius / 100 in the same unit as Petrosian radius (here, ).
-Using the pixel size we calculate the angular size of galaxies in radian.
+Using the default SDSS pixel size of 0.396 arcsec/pixel, we calculate the angular size of galaxies in radian.
 Then we obtained the physical sizes of galaxies by multiplying angular sizes and angular distances.
@@
 3).
+Since the identification of BP/X structures is affected by the inclination of galaxies, it is worth discussing the inclination distribution of our sample.
+We measure the inclination angle using the axis ratio obtained from SDSS data, applying the relation (Holmberg, 1958), where is the intrinsic axial ratio.
+Following Tully & Fisher (1977), we adopt for this analysis.
+We find a median inclination of degree, with a – range of – degrees.
+For an ideally edge-on galaxy, the inclination angle is degree, indicating that our sample largely consists of nearly edge-on galaxies.
 7 Result
@@
 First, we investigate the differences in properties like mass, disk radius, disk height and axial ratio among the different classes we have identified in the whole sample.
-Fig.
-4, we plot the distributions of mass, axis ratio, disk height, and disk radius for for samples in this study.
+In Fig.
+4, we plot the distributions of mass, axis ratio, disk height, and disk radius for the four samples in this study.
 The control, both-sided, one-sided, and strong both-sided are respectively shown with blue, green, red, and purple colors.
-To better understand the distribution, we fitted each dataset using the ’log normal’ distribution.
-For fitting, we used the Python package scipy .777scipy
-The distribution of mass (4a) shows that the mass distribution for the samples without any signature of bulge formation (control sample) lies in a range of .
+To better understand the distribution, we fitted each dataset using the ‘log normal’ distribution.
+For fitting, we used the Python package Scipy.
+The distribution of masses (4a) shows that the mass distribution for the sample without any signature of bulge formation (control sample) peaks at .
 But for the other 3 samples, i.e.
-the ones showing signatures of bulge formation goes up to nearly .
-This is expected because bulge formation occurs in galaxies having a bar and high mass galaxies are more likely to have bars than the lower mass galaxies (Erwin (2018)).
-We have defined axial ratio as the ratio of disk height to disk radius.
-For the galaxies having signature of bulge formation, we expect the axial ratio to be higher than that of the control sample as the disk height is higher for them.
-For our lognorm fit, we obtained the mean value of axial ratio of strong both-sided galaxies is higher among all samples (4b).
-For all our further analysis described in section 7.4 and 7.5, we have used only the galaxies having mass, disk radius and disk height less than , 14kpc and 5kpc respectively.
-By using these limits, we were able to remove the relatively higher values in the data so that we could study the remaining data in a more detailed manner.
-These limits do not affect the correctness of the analysis because this includes nearly 98.8% of the whole galaxies.
-Category-wise, we have more than 99% of control sample and one-sided sample , nearly 98.3% of both-sided and 97% of strong both-sided inside this limit.
+the ones showing signatures of bulge formation peaks in the range of .
+This is expected because bulge formation occurs in galaxies having a bar and high mass galaxies are more likely to host bars than the lower mass galaxies (Erwin (2018)).
+We have defined axial ratio as the ratio of disk scale height to disk scale radius.
+For the galaxies having signature of bulge formation, we expect the axial ratio to be higher than that of the control sample as bars thicken during BP/X bulge formation.
+From our ‘lognorm fit’, we obtained the mean value of axial ratio of strong both-sided galaxies higher among all samples (4b), though not significantly different.
+This similarity in the axial ratio is also evident from the scale height (4c) and scale radius distributions (4d).
+For further analysis described in Section 7.4 and 7.5, we have used only the galaxies having stellar mass, disk radius and disk height less than , 14 kpc, and 5 kpc, respectively.
+By using these limits, we remove relatively higher value outliers in the data.
+These limits do not affect the correctness of the analysis because the sample includes nearly 98.3% of galaxies.
+Category-wise, we have nearly 99% of control sample and one-sided sample, nearly 97.4% of both-sided and 95.4% of strong both-sided within this limit.
 7.2 BP/X fraction
@@
 Lütticke et al.
-(2000) presented the first detailed statistics of BP/X bulges using a large sample of edge-on galaxies from Third Reference Catalogue of Bright Galaxies (RC3: de Vaucouleurs et al., 1991).
+(2000) presented the first detailed statistics of BP/X bulges using a large sample of edge-on galaxies from Third Reference Catalog of Bright Galaxies (RC3: de Vaucouleurs et al., 1991).
 They found BP/X bulges in their sample of 734 classifiable bulges.
@@
 In their clean sample of 1329 galaxies in band, they reported BP/X structures.
-Since Yoshino & Yamauchi (2015) used SDSS sample, we cross-matched our sample with the their catalog of 1312 edge-on galaxies in the -band, and identified 439 galaxies in common.
+Since Yoshino & Yamauchi (2015) used SDSS sample, we cross-matched our sample with the their catalog of 1312 edge-on galaxies in the -band, and identified 455 galaxies in common.
 The primary reason for the discrepancy in sample sizes lies in the differing selection criteria.Yoshino & Yamauchi (2015) selected galaxies with a minor-to-major isophotal axis ratio in the -band less than 0.25, whereas our sample was selected based on volunteer votes.
 When we compare the values from their model fitting for both the common and complete samples, we find that the common sample predominantly consists of galaxies with , while complete sample show median , indicating that a large fraction of their sample comprises moderately inclined systems.
-Nonetheless, out of these common galaxies they identified 123 total BP/X structures (7 strong, 40 standard, and 76 weak in their definitions) and we identified 150 total BP/X structures (65 strong both-sided and 85 both-sided).
+Nonetheless, out of these common galaxies they identified 129 total BP/X structures (7 strong, 42 standard, and 80 weak in their definitions) and we identified 165 total BP/X structures (68 strong both-sided and 97 both-sided).
 Recently, a similar fraction of BP/X bulges is reported in -band images.
@@
 As described previously in Section 4, we started with 6653 edge-on galaxies and classified them into three classes of BP/X bulges: Both-sided, Strong both-sided, and One-sided.
-We identify 1675 () both-sided bulges, and 1108 () one-sided bulges.
-Among both-sided bulges 545 () are strong both-sided with prominent X-structure, and remaining 1130 () show faint X-structure.
-When combined together, both-sided and one-sided make (2783 galaxies) of the total sample.
+We identify 1673 () both-sided bulges, and 1112 () one-sided bulges.
+Among both-sided bulges 504 () are strong both-sided with prominent X-structure, and remaining 1169 () show faint X-structure.
+When combined together, both-sided and one-sided make (2785 galaxies) of the total sample.
 Though our base sample is nearly half of that studied in Marchuk et al.
@@
 In Fig.
-4(a), we have plotted the stellar mass distribution of both-sided (gold), strong both-sided (green), one-sided (black), and control sample (cyan) in this study.
+4a, we have plotted the stellar mass distribution of both-sided (green), strong both-sided (purple), one-sided (red), and control sample (blue) in this study.
 We found that the distribution of control galaxies peaks at lower stellar mass than the galaxies with BP/X bulges.
@@
 Fig.
-5 shows the fraction of both-sided (green), strong both-sided (orange), one-sided (blue) bulges with the stellar mass of galaxies.
+5 shows the fraction of both-sided (green), strong both-sided (purple), one-sided (red) bulges with the stellar mass of galaxies.
 The vertical bars in respective curve represent statistical uncertainty in measurement defined as , where being fraction of a bulge class at any stellar mass bin and being total number of galaxies in that bin.
@@
 Additionally, the fraction of both-sided BP/X bulges, which also includes strong both-sided bulges, is always the highest for any stellar mass.
-In the shown stellar mass range, we can see an increase in strong both-sided, one-sided and both-sided bulge fraction, respectively, from 0, 0, and 12 to 18, 24, and 38 percent.
+In the shown stellar mass range, we can see an increase in strong both-sided, one-sided and both-sided bulge fraction, respectively, from 0, 0, and 20 to 18, 26, and 42 percent.
 7.4 Dependence of BP/X bulge on stellar surface density
@@
 The low surface density disks () show nearly constant fraction of both-sided and strong both-sided BP/X bulges.
-Below this surface density limit, the fraction of strong both-sided structures is less than , while both-sided makes about , and the contribution of one-side structures rises from zero to .
+Below this surface density limit, the fraction of strong both-sided structures is less than , while both-sided makes about , and the contribution of one-side structures is about .
 For high surface density disks (), the fraction of BP/X structure hosting disks rises by about in all three groups.
@@
 Followings are key summarizing points of our study.
-We visually classified 6653 edge-on galaxies into 1675 () both-sided bulges, and 1108 () one-sided bulges.
-Among both-sided bulges 545 () are strong both-sided with prominent X-structure, and remaining 1130 () show faint X-structure.
-When combined together, both-sided and one-sided make (2783 galaxies) of the total sample.
-Our sample provides the largest catalog of BP/X bulges to date, which includes a category of one-sided BP/X bulges.
+We visually classified 6653 edge-on galaxies into 1673 () both-sided bulges, and 1112 () one-sided bulges.
+Among both-sided bulges 504 () are strong both-sided with prominent X-structure, and remaining 1169 () show faint X-structure.
+When combined together, both-sided and one-sided make (2785 galaxies) of the total sample.
+Our sample provides the largest catalog of BP/X bulges to date, which includes a category of candidate one-sided BP/X bulges.
 The catalog of one-sided structures is one of its kind sample, which likely contains samples of ongoing buckling events in edge-on galaxies.
@@
 The fraction of both-sided BP/X bulges is always the highest for any stellar mass.
-At , the fraction of strong both-sided, one-sided and both-sided bulge is 18, 24, and 38 percent, respectively.
+At , the fraction of strong both-sided, one-sided and both-sided bulge is 18, 26, and 42 percent, respectively.
 We report a strong dependence of BP/X fraction on average stellar surface density of galaxy.
@@
 Future studies employing detailed 2D decomposition for scale size measurements and SED fitting for stellar mass, star formation rate, dust content, and star formation history will be instrumental in the era of large-scale surveys, e.g., Euclid, LSST, 4MOST, and Roman.
+With the increase in spatial resolution, it will be possible to distinguish between classical bulge and weak BP/X bulge.
+We can find the real one-sided BP/X bulge galaxies from the class of one-sided galaxies.
 Additionally, our sample can be used to train AI models for studying large datasets.
@@
 - Di Matteo et al. (2019) Di Matteo, P., Haywood, M., Lehnert, M., et al. 2019, Astronomy & Astrophysics, 632, A4
-- D´ıaz-Garc´ıa et al. (2020) Díaz-García, S., Moyano, F. D., Comerón, S., et al. 2020, A&A, 644, A38
-- D´ıaz-Garc´ıa et al. (2016) Díaz-García, S., Salo, H., Laurikainen, E., & Herrera-Endoqui, M. 2016, Astronomy & Astrophysics, 587, A160
+- Díaz-García et al. (2020) Díaz-García, S., Moyano, F. D., Comerón, S., et al. 2020, A&A, 644, A38
+- Díaz-García et al. (2016) Díaz-García, S., Salo, H., Laurikainen, E., & Herrera-Endoqui, M. 2016, Astronomy & Astrophysics, 587, A160
 - Donohoe-Keyes et al. (2019) Donohoe-Keyes, C., Martig, M., James, P., & Kraljic, K. 2019, Monthly Notices of the Royal Astronomical Society, 489, 4992
@@
 - Galloway et al. (2015) Galloway, M. A., Willett, K. W., Fortson, L. F., et al. 2015, MNRAS, 448, 3442
+- Garland et al. (2024) Garland, I. L., Walmsley, M., Silcock, M. S., et al. 2024, Monthly Notices of the Royal Astronomical Society, 532, 2320
 - Ghosh et al. (2024) Ghosh, S., Fragkoudi, F., Di Matteo, P., & Saha, K. 2024, A&A, 683, A196
 - Ginsburg et al. (2019) Ginsburg, A., Sipőcz, B. M., Brasseur, C. E., et al. 2019, AJ, 157, 98
-- Gonzalez et al. (2017) Gonzalez, O. A., Debattista, V. P., Ness, M., Erwin, P., & Gadotti, D. A. 2017, Monthly Notices of the Royal Astronomical Society: Letters, 466, L93
 - Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357–362
 - Heller & Shlosman (1994) Heller, C. H. & Shlosman, I. 1994, Astrophysical Journal, Part 1 (ISSN 0004-637X), vol. 424, no. 1, p. 84-105, 424, 84
+- Holmberg (1958) Holmberg, E. 1958, Meddelanden fran Lunds Astronomiska Observatorium Serie II, 136, 1
 - Hunter (2007) Hunter, J. D. 2007, Computing in Science Engineering, 9, 90
@@
 - Marchuk et al. (2022) Marchuk, A. A., Smirnov, A. A., Sotnikova, N. Y., et al. 2022, MNRAS, 512, 1371
-- Marchuk et al. (2022) Marchuk, A. A., Smirnov, A. A., Sotnikova, N. Y., et al. 2022, Monthly Notices of the Royal Astronomical Society, 512, 1371
 - Martin & Friedli (1997) Martin, P. & Friedli, D. 1997, Astronomy and Astrophysics, v. 326, p. 449-464, 326, 449
@@
 - Stoughton et al. (2002) Stoughton, C., Lupton, R. H., Bernardi, M., et al. 2002, The Astronomical Journal, 123, 485
-- Stoughton et al. (2002) Stoughton, C., Lupton, R. H., Bernardi, M., et al. 2002, AJ, 123, 485
 - Toomre (1981) Toomre, A. 1981, in Structure and Evolution of Normal Galaxies, ed. S. M. Fall & D. Lynden-Bell, 111–136
 - Trujillo et al. (2006) Trujillo, I., Förster Schreiber, N. M., Rudnick, G., et al. 2006, ApJ, 650, 18
+- Tully & Fisher (1977) Tully, R. B. & Fisher, J. R. 1977, A&A, 54, 661
 - Van der Kruit & Searle (1981) Van der Kruit, P. & Searle, L. 1981, Astronomy and Astrophysics, vol. 95, no. 1, Feb. 1981, p. 105-115., 95, 105